Please, help me to simplify the expression: 4a⁸/2a⁴
*Using only positive exponents

Answer by stanbon(75887)

simplify the expression: 4a⁸/2a⁴
-------------------
= (4/2)(a^8/a^4) = 2a^(8-4) = 2a^4
-----------------
Cheers,
Stan H.
-----------------